OpenStudy (anonymous):

The sequence an = 5n + 3 is the same as the sequence a1 = 8, an = an-1 + 5. TRUE or FALSE **Please explain! :) Thanks! :)

OpenStudy (anonymous):

the 1st equation is general,2nd is recursive,first write the sequence

OpenStudy (anonymous):

5(1)+3=8 5(2)+3=13 5(4)+3=23

OpenStudy (anonymous):

oh idk how to haha i'm like super confused :(

OpenStudy (anonymous):

how come u skipped 5(3)=15?

jimthompson5910 (jim_thompson5910):

an = an-1 + 5 basically says "to get any term, take the previous term and add 5 to it"

OpenStudy (anonymous):

oops i meant 5(3)+3=18

OpenStudy (anonymous):

hi @jim_thompson5910 thanks for coming! :) so does that mean this statement is false?? because the first one says +3 not +5?

jimthompson5910 (jim_thompson5910):

so a1 = 8, an = an-1 + 5 means "start with 8, then add 5 to each previous term to get the next term to generate the sequence"

OpenStudy (anonymous):

8,13,18...5(n)+3 the ? is is this the same as \[a _{n}=a _{n-1}+5\]

OpenStudy (anonymous):

as jim_thompson5910 explained

OpenStudy (anonymous):

i don't think so right? cuz the first one is +3 and the second one is +5 right?? SO this statement is FALSE?

jimthompson5910 (jim_thompson5910):

If you start an = 5n + 3 with n = 1, then a1 = 5(1) + 3 = 8 ----> a1 = 8 a2 = 5(2) + 3 = 13 ----> a2 = 13 a3 = 5(3) + 3 = 18 ----> a3 = 18 a4 = 5(4) + 3 = 23 ----> a4 = 23 ------------------------------------------------------- Compare this to the rule "start with 8, then add 5 to each previous term to get the next term to generate the sequence" First term: 8 Second term: 8+5 = 13 Third term: 13+5 = 18 Fourth term: 18+5 = 23 So it's the same sequence

OpenStudy (anonymous):

oh it is?? is it the same sequence if the outcomes are the same right?? it doesn't matter if it is +3, +5 etc,,, as long as the terms are the same (like you have showed?)

OpenStudy (anonymous):

So this statement is TRUE?

jimthompson5910 (jim_thompson5910):

yes, your logic and line of thinking on that is correct

jimthompson5910 (jim_thompson5910):

Those are just 2 ways to say the same thing (which is 8, 13, 18, 23, ...)

OpenStudy (anonymous):

so final answer is TRUE because of that?

OpenStudy (anonymous):

yes its true cos you get the terms by adding 5

OpenStudy (anonymous):

Cool!! THanks for the help everybody!!!! =) =D ;) :D

OpenStudy (anonymous):

i totally understand it now!! :)